OpenStudy (anonymous):

The circle given by x^2+y^2-4x-10=0 can be written in standard form like this: (x-h)^2+y^2=14. What is the value of h in this equation??

OpenStudy (anonymous):

lets rearrange it a bit x^2-4x+y^2=10 Do you know how to complete the square?

OpenStudy (anonymous):

No :( can you help me?

OpenStudy (nincompoop):

there should be a very very easy way to figure this out

OpenStudy (campbell_st):

you need to complete the square in x and the y's are already a perfect square so what needs to be added to both sides to to have a perfect square in x..?

OpenStudy (anonymous):

would you add 4x to both sides?

OpenStudy (nincompoop):

>.<

OpenStudy (anonymous):

no?

OpenStudy (ranga):

x^2+y^2-4x-10=0 group x terms together and y terms together: x^2 - 4x + y^2 - 10 = 0 complete the square for x and y separately. There is a y^2 but no y and therefore y is already a perfect square. complete the square for x^2 - 4x divide the coefficient of x by 2: -4/2 = -2. The -2 will go inside the square. Subtract (-2)^2 x^2 - 4x = (x - 2)^2 - 4 y^2 = (y - 0)^2 Put them back in the original equation. Take the constant to the right.

OpenStudy (anonymous):

would h, the constant, be 2?

OpenStudy (ranga):

yes. h = 2 and k = 0

OpenStudy (anonymous):

Thank you :) for this problem do you know what the value of b would be for this circle in general form? x^2+y^2+ax+by+c=0

OpenStudy (ranga):

In the diagram, first locate the center of the circle and determine its coordinates. Also find the radius. What are they?

OpenStudy (ranga):

If you are having difficulty quickly locating the center of the circle you can do it the slightly long way. What is the maximum y value the circle attains? What is the minimum y value the circle attains. Add them up and divide by 2 and that will be the y coordinate of the center. Do the same with x. (Max x-value + Min. x value) / 2

OpenStudy (anonymous):

coordinates are (1,-2) and radius is 3 correct?

OpenStudy (ranga):

Good job! Now plug that into the standard form for the equation of a circle: (x - h)^2 + (y - k)^2 = r^2 where (h, k) is the center and r is the radius.

OpenStudy (anonymous):

(x-1)^2+(y- -2)^2=9

OpenStudy (anonymous):

how do i know what the value of b is though?

OpenStudy (ranga):

We will get there. Write (y --2) as (y + 2) in the general equation first.

OpenStudy (anonymous):

okay so (x-1)^2+(y+2)^2=9

OpenStudy (ranga):

Now expand the terms. Here they are interested only in b and not a and c and so it would be sufficient to expand just the (y+2)^2. But for practice expand (x-1)^2 also.

OpenStudy (ranga):

For example, (x-1)^2 = x^2 - 2x + 1

OpenStudy (anonymous):

would that be (x-1)(x+1)+(y-2)(y-2)

OpenStudy (ranga):

yes. Or, if you remember the identity (a - b)^2 = a^2 - 2ab + b^2 you can use that too. Or you can multiply (a+b)(a+b) if you don't know the identity.

OpenStudy (anonymous):

so b would be 1?

OpenStudy (anonymous):

wouldn't that be (y-2)(y+2) ?

OpenStudy (ranga):

No. You have to expand (y+2)^2 first.

OpenStudy (ranga):

(y+2)(y+2) = ?

OpenStudy (anonymous):

y^2+4?

OpenStudy (ranga):

No. (y + 2)(y + 2) = y^2 + 2y + 2y + 4 = y^2 + 4y + 4

OpenStudy (anonymous):

oh okay, so b would be 4?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

thank you :)

OpenStudy (ranga):

Had they asked you to find a, b and c, then you will expand the whole thing: (x-1)^2+(y+2)^2=9 x^2 - 2x + 1 + y^2 + 4y + 4 = 9 x^2 + y^2 - 2x + 4y - 4 = 0 Now you can compare it to the general form of the equation: x^2 + y^2 + ax + by + c = 0 and conclude a = -2; b = 4 and c = -4

OpenStudy (ranga):

yw.